OpenStudy (anonymous):

how do u completely factorize 5z squared - 5?

OpenStudy (anonymous):

Difference of squares

OpenStudy (anonymous):

\[ 5z^2-5 = (\sqrt{5}z-\sqrt{5})(\sqrt{5}z+\sqrt{5}) \]

OpenStudy (anonymous):